# Hanging on the cliff: EMRI formation with local two-body relaxation and post-Newtonian dynamics

### v1

@@ 3.1 Numerical overview and setup
In this work we employ an extended version of the code presented in Bonetti et al. (2016), a PN few-body integrator with adaptive step size, which we use to study the evolution of a binary system comprised of an MBH and a stellar-mass BH. The integrator leverages on a C++ implementation of the Bulirsch-Stoer algorithm (Bulirsch & Stoer 1966) coupled with Richardson extrapolation (Richardson 1911) to increase the accuracy of the numerical solution. The adaptive time step allows to get smaller steps when the dynamics is faster, such as at the pericentre of the orbit.
@@
Here, $H_{0}$ is the non-relativistic Newtonian Hamiltonian, while $H_{1}$, $H_{2}$ and $H_{2.5}$ are corrective terms introducing purely relativistic effects. Complete expressions for these terms, matching the form implemented numerically, can be found for example in Bonetti et al. (2016) [^7] Bonetti et al. (2016) focus on the three-body problem, and therefore reports three-body PN Hamiltonian terms. We consistently use the same Hamiltonian, and reduce the evolution to a two-body problem by placing a third body with negligible mass at very large distance from the inner binary, effectively isolating its dynamics.. $H_{2.5}$ is the first non-conservative term and accounts for GW radiation (Jaranowski & Schäfer 1997; Blanchet 2014). We only consider Schwarzschild BHs, thus we ignore the 0.5PN and 1.5PN terms of the expansion, as they account for spin and spin-orbit effects (Schäfer & Jaranowski 2024).
@@
At the same time we forcefully keep the MBH fixed and still at the origin of our Cartesian coordinate system. This choice is motivated by the fact that in a nuclear cluster the MBH will simultaneously interact with virtually all the objects in the stellar distribution, which will only cause a small Brownian motion-like displacement from the center. We ignore such stochastic displacement of the MBH, even if it might be relevant as the number of stellar objects and the mass of the MBH gets lower (Merritt 2013a). Together with the evolution of the binary system, the code can account for the presence of additional external gravitational potentials, possibly representing a distribution of background bodies. We stress again that only a single stellar-mass BH is evolved by directly solving Hamilton’s equations.
Figure 5: Evolution of a stellar-mass BH over few passages of the apocentre, using our local treatment of two-body relaxation. The top right panel shows the 3D trajectory of the stellar-mass BH, while the other three panels show its projections onto the three planes of the reference system. The colour of the curve shows the time elapsed since the beginning of the evolution. Red dots indicate the points were a velocity kick is given and they coincide with the time step of the main integrator. Points in between red dots are interpolated. The black cross shows the initial position of the stellar-mass BH, while the MBH is fixed at $(0,0,0)$. Note that the $y$ and $z$ axes have different scales for better readability. In this example, $M_{\bullet}=4\times 10^{6}\,\mathrm{M}_{\sun}$, $a_{0}=0.1\,\mathrm{pc}$, $e_{0}=0.999$.
@@
    before moving to the next time step, we update the momentum of the stellar-mass BH to ${p^{\prime}}=\@vec{p}+m_{\bullet}\Delta\@vec{v}$, where $m_{\bullet}$ is the mass of the stellar-mass BH [^8] Here we use the relation ${p}=m\@vec{v}$, which is not valid at PN order. Thus, while the evolution of the system is PN, the small velocity kicks we give to the stellar-mass BH at the end of each time step are Newtonian. .
Following this procedure, we can statistically account for two-body encounters locally and multiple times within a single orbit, without leveraging on orbit-by-orbit approaches (QS23) or orbit-averages of the diffusion coefficients (Zhang & Amaro Seoane 2024).
@@ 3.2 Local treatment of two-body relaxation
As explained in appendix A, in order to compute local diffusion coefficients we have to first find $f_{\alpha}(\mathcal{E})$ given $n_{\alpha}$ and then integrate it over $\mathcal{E}$. This process involves nested integrals and derivatives, making unpractical to numerically perform such computations on the fly at each time step of our simulations. Thus, we precompute the auxiliary functions $\mathcal{I}_{k,\alpha}$ and $\mathcal{J}_{k,\alpha}$, that are directly related to the diffusion coefficients, for some logarithmically spaced values of $\mathcal{E}$ and $\psi$, and then, during the execution, we bilinearly interpolate these values of $\mathcal{I}_{k,\alpha}$ and $\mathcal{J}_{k,\alpha}$. In this way, we get approximate local values for those coefficients given the exact position and velocity of the stellar-mass BH at a particular time step.
@@
In these equations, $\Delta t$ is the length of the particular time step we are considering. Thus, the values of $\Delta v_{\parallel}$ and $\Delta v_{\perp}$ we draw represent a statistical realisation of the change in velocity due to all of the encounters that have occourred during that time step.
@@ 3.3 Orbit-averaged treatment of two-body relaxation
Figure 6: Sketch depicting an example orbit change following our orbit-averaged procedure. The stellar-mass BH, initially on the orbit labeled as 00, is kicked when $r=a$, and ends on the orbit labeled as $1$. Both orbits have the same semi-major axis $a$, but different eccentricities. The dashed circle represents a circular orbit with radius $a$. In this example, there is no way to rotate either orbit around the focus occupied by the MBH such that the pericentre or apocentre of orbit 0 crosses at any point orbit 1. Thus, neither pericentres nor apocentres are good spots to employ our orbit-averaged procedure. Note that the eccentricities of these example orbits are much lower compared to the ones investigated in this work. Moreover, the difference between subsequent orbits is much exaggerated here.
Figure 7: Evolution of a stellar-mass BHs over few orbits with the orbit-averaged treatment of two-body relaxation. The top right panel shows the 3D trajectory of the stellar-mass BH, while the other three panels show its projections onto the three planes of the reference system. The colour of the curve shows the time elapsed since the beginning of the evolution. Red dots indicate the points were a velocity kick is given, close to $r=a$. We let the body complete the first orbit without applying the kicks to visualise its trajectory better. The black cross shows the initial position of the stellar-mass BH, while the MBH is fixed at $(0,0,0)$. Note that the $y$ and $z$ axes have different scales for better readability. In this example, $M_{\bullet}=4\times 10^{6}\,\mathrm{M}_{\sun}$, $a_{0}=0.1\,\mathrm{pc}$, $e_{0}=0.997$.
In appendix B, we show that, once $\delta v_{\parallel}$ and $\delta v_{\perp}$ are drawn, then
@@
In appendix C, we show that a suitable orbit-integrated kick in velocity can be computed from $\Delta E$ and $\Delta J$ as:
@@
When updating the orbit, we change the velocity of the stellar-mass BH without displacing it. This is possible only if the old and the new orbit intersect at the location where $\Delta v$ is applied, a fact that we ensure by applying the kick when $r\simeq a$ (Fig. 6 illustrates the situation with a sketch). It is safe to assume that the orbits intersect there, as long as the energy of the particle (and thus $a$, since $\mathcal{E}\simeq GM_{\bullet}/2a$) does not change drastically along a single period [^10] We stress that the $\Delta\mathcal{E}$ and $\Delta J$ we are using here only come from the effects of two-body relaxation. GWs emission can strongly and quickly modify the energy of the orbit, but their effect is only included in the 2.5PN term of the Hamiltonian.. In case this assumption is not verified, the orbit-averaging approximation is not valid anymore, as the time needed for two-body relaxation to significantly change $\mathcal{E}$ is less than the orbital period (Binney & Tremaine 2008; Merritt 2013a).
@@ 3.4 Stopping criteria and initial conditions
As we aim at gathering a statistically significant number of EMRIs and DPs, we choose to limit the parameter space that we explore to a sub-portion close to the loss cone edge. Moreover, as we assume a steady-state stellar profile, we are interested in exploring orbital parameters for which the relaxation time $t_{\mathrm{rlx}}$ is not too large, as the initial conditions and the diffusion coefficients would be affected by the overall relaxation of the system. In particular, we consider as our timescale of investigation ${\sim}1\%$ of the time-to-peak of (classical) EMRI rates in the systems we are considering, which estimates the timescale of evolution of the distribution function of scatterers (Broggi et al. 2022):
@@
where we leverage on the the M-$\sigma$ relation (Gültekin et al. 2009) to obtain the final scaling with the MBH mass.
We initialise all our simulations inside the region of interest where $t_{\mathrm{rlx}}<t_{\mathrm{max}}$ following a procedure described at the end of this section and we stop and discard any run where the orbit leaves such region. This choice does not affect the EMRI-to-DP ratio. In fact, $t_{\mathrm{max}}$ is so long with respect to the orbital period that any plunge or EMRI must enter this region before the capture. As long as the procedure is not biased towards EMRI or DP candidates, once a particle is discarded we are safe to reinitialise it as an uncorrelated particle.
The reinitialisation procedure becomes biased when the restriction to $t_{\mathrm{rlx}}<t_{\mathrm{max}}$ excludes a significant part of the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ curve. In this case, it can happen that some orbits well inside the GW emission dominated region but still far from the loss cone, can reach the $t_{\mathrm{rlx}}=t_{\mathrm{max}}$ boundary before crossing the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ separatrix, thus being erroneously stopped and reinitialised instead of being counted as EMRIs. A clear example of such orbit is shown by the blue track in Fig. 8: the event will clearly produce an EMRI, but according to our reinitialisation criterion is effectively discarded and resimulated. As DP candidates are not affected by this issue, the EMRI-to-DP ratio is influenced by these discarded EMRIs. This problem mostly affects the EMRI-to-DP ratio at small initial semi-major axes for large $M_{\bullet}$, as we show in Sect. 4.1.
For this reason, we correct our counts of EMRIs to also include orbits which cross into the excluded region while inside the GW-dominated region. Before accepting the orbit as an EMRI, we also verify that the final value of $a$ is below the evolution track of an orbit which evolves as Eq. (32), where $c_{0}$ is chosen so that such track crosses the minimum of the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ curve (again, see Fig. 8). This means that, ignoring stochastic kicks due to two-body encounters, the orbit would eventually reach our condition for EMRI formation, while inside the excluded region.
Figure 8: Depiction of two example orbit evolutions on the $(1-e,a)$ plane. The blue track shows the formation of an EMRI which is missed by our standard definition, and requires the correction described in the text to be rightfully counted. The pink track instead shows a regular reinitialisation. We also plot in dashed brown the evolution track of an orbit which passes through the minimum of the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ curve following Eq. (32). It can be seen that all orbits which cross into the excluded region while being in the GW-dominated region fall below this evolution track, meaning they will reach the solid green curve, as long as the stochastic kicks due to two-body relaxation are small. See Fig. 1 for the meaning of the other elements. Here we set $M_{\bullet}=4\times 10^{6}\,\mathrm{M}_{\sun}$.
@@
at the corresponding energy $\mathcal{E}_{0}=GM_{\bullet}/2a_{0}$. Then, we randomly draw an initial value for $\mathcal{R}_{0}$ from the probability distribution functon
@@ 4 Results
In this section, we present the results we obtained running a large number of simulations with the code described in Sect. 3.
@@ 4.1 EMRI-to-DP ratio and cliffhanger EMRIs
Figure 10: Distribution of the maximum ratio between $a$ evaluated at subsequent apocentres for EMRIs for selected values of $M_{\bullet}$ and $a_{0}$. The black dotted vertical line shows $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}=2$, which roughly delimits cliffhanger EMRIs to its right. In the left panel, the orange dashed and green dashdotted vertical lines show $a_{0}/a_{\mathrm{f}}^{\mathrm{lc}}$ from Eq. (38), which is a rough prediction of the largest possible $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}$ ratio. Each histogram is normalised to 1. Note the different $x$-axis scale as well as binning strategy in the two panels.
@@
We report our main findings in the left panel of Fig. 9. For $M_{\bullet}=10^{6}\,\mathrm{M}_{\sun}$ and $M_{\bullet}=4\times 10^{6}\,\mathrm{M}_{\sun}$, we recover the expected dichotomy between EMRIs and DPs. In this case, EMRIs can only form from orbits with $a_{0}\lesssim a_{\mathrm{c}}$, while those with $a_{0}\gtrsim a_{\mathrm{c}}$ exclusively result in DPs. We find that the transition between the two regimes occurs over about an order of magnitude in $a_{0}$ at $a_{\mathrm{c}}\simeq 10^{-2}R_{\mathrm{inf}}$, which is consistent within a factor of two with previous works (Hopman & Alexander 2005; Raveh & Perets 2021; Qunbar & Stone 2023). We believe that the quantitative difference is driven by differences in the nuclear cluster models, in the accuracy of the dynamical model (e.g. PN order) and in the treatment of two-body relaxation. We better explore the impact that some of these factors have on the shape of $S(a_{0})$ in Sect. 4.2 and 4.3.
@@
Conversely, for $M_{\bullet}<10^{6}\mathrm{M}_{\sun}$, we observe a relevant number of EMRIs formed from initially wide orbits, with $a_{0}\gg a_{\mathrm{c}}$. These events are consistent with the cliffhanger EMRI population first seen in Monte Carlo simulations by QS23 and, to the best of our knowledge, never found in previous work. We observe cliffhanger EMRIs forming for $M_{\bullet}\lesssim 3\times 10^{5}\,\mathrm{M}_{\sun}$ and they appear to be more common as $M_{\bullet}$ decreases, consistently with the argument based on the $(1-e,a)$ phase space presented in Sect. 2.4.
@@
Our results for $S(a_{0})$ are compared to those obtained by QS23 and other authors in the right panel of Fig. 9. First, we can notice the large scatter in $a_{\mathrm{c}}/R_{\mathrm{inf}}$ that spans almost an order of magnitude, from $\approx 3\times 10^{-3}$ (QS23) to $\approx 2\times 10^{-2}$ (Raveh & Perets 2021). This is likely due to the differences in the parameters of the simulated systems as well as in the treatment of the dynamics and loss cone definition, as we will discuss in Sect. 4.3. Focusing on cliffhanger EMRIs, QS23 observe their formation only for MBHs with $M_{\bullet}\lesssim 5\times 10^{4}\mathrm{M}_{\sun}$, a threshold set almost an order of magnitude below our limiting value. Moreover, we see a much larger fraction of EMRIs over DPs at large $a_{0}$, with a maximum ratio of $S(a_{0})\simeq 0.65$ at $a_{0}\gg 10^{-2}R_{\mathrm{inf}}$ for $M_{\bullet}=10^{4}\mathrm{M}_{\sun}$. As a further difference, our $S(a_{0})$ functions hint at a downward trend for $a_{0}\gg 10^{-2}R_{\mathrm{inf}}$, as opposed to tending to a constant at $a_{0}\to\infty$. This could be due to the fact that QS23 stopped their investigation at $a_{\mathrm{max}}\simeq R_{\mathrm{inf}}/3$, as their implementation does not account for the stellar potential.

### v2

@@ 3.1 Numerical overview and setup
Figure 5: Evolution of the position of a stellar-mass BH over few passages of the apocentre, using our local treatment of two-body relaxation. The top right panel shows the 3D trajectory of the stellar-mass BH, while the other three panels show its projections onto the three planes of the reference system. The colour of the curve shows the time elapsed since the beginning of the evolution. The red dots indicate the points where a velocity kick is given and they coincide with the time step of the main integrator. The points between the red dots are interpolated. The black cross shows the initial position of the stellar-mass BH, while the MBH is fixed at $(0,0,0)$. We note that the $y$- and $z$-axes have different scales for better readability. In this example $M_{\bullet}=4\times 10^{6}\,\mathrm{M}_{\sun}$, $a_{0}=0.1\,\mathrm{pc}$, $e_{0}=0.999$.
In this work we employ an extended version of the code presented in Bonetti et al. (2016), a PN few-body integrator with adaptive step size, which we use to study the evolution of a binary system comprised of an MBH and a stellar-mass BH. The integrator leverages on a C++ implementation of the Bulirsch-Stoer algorithm (Bulirsch & Stoer 1966) coupled with Richardson extrapolation (Richardson 1911) to increase the accuracy of the numerical solution. The adaptive time step allows us to get smaller steps when the dynamics is faster, such as at the pericentre of the orbit.
@@
Here, $H_{0}$ is the non-relativistic Newtonian Hamiltonian, while $H_{1}$, $H_{2}$ and $H_{2.5}$ are corrective terms introducing purely relativistic effects. Complete expressions for these terms, matching the form implemented numerically, can be found for example in Bonetti et al. (2016). [^7] Bonetti et al. (2016) focus on the three-body problem, and therefore report three-body PN Hamiltonian terms. We consistently use the same Hamiltonian, and reduce the evolution to a two-body problem by placing a third body with negligible mass at very large distance from the inner binary, effectively isolating its dynamics. $H_{2.5}$ is the first non-conservative term and accounts for GW radiation (Jaranowski & Schäfer 1997; Blanchet 2014). We only consider Schwarzschild BHs, thus we ignore the 0.5PN and 1.5PN terms of the expansion, as they account for spin and spin-orbit effects (Schäfer & Jaranowski 2024).
@@
At the same time we forcefully keep the MBH fixed and still at the origin of our Cartesian coordinate system. This choice is motivated by the fact that in a nuclear cluster the MBH will simultaneously interact with virtually all the objects in the stellar distribution, which will only cause a small Brownian motion-like displacement from the centre. We ignore such a stochastic displacement of the MBH, even if it might be relevant as the number of stellar objects and the mass of the MBH gets lower (Merritt 2013a). Together with the evolution of the binary system, the code can account for the presence of additional external gravitational potentials, possibly representing a distribution of background bodies. We note again that only a single stellar-mass BH is evolved by directly solving Hamilton’s equations.
@@
    before moving to the next time step, we update the momentum of the stellar-mass BH to ${p^{\prime}}=\@vec{p}+m_{\bullet}\Delta\@vec{v}$, where $m_{\bullet}$ is the mass of the stellar-mass BH. [^8] Here we use the relation ${p}=m\@vec{v}$, which is not valid at PN order. Thus, while the evolution of the system is PN, the small velocity kicks we give to the stellar-mass BH at the end of each time step are Newtonian.
Following this procedure, we can statistically account for two-body encounters locally and multiple times within a single orbit, without leveraging on orbit-by-orbit approaches (QS24) or orbit-averages of the diffusion coefficients (Zhang & Amaro Seoane 2024).
@@ 3.2 Local treatment of two-body relaxation
As explained in Appendix A, in order to compute local diffusion coefficients we have to first find $f_{\alpha}(\mathcal{E})$ given $n_{\alpha}$ and then integrate it over $\mathcal{E}$. This process involves nested integrals and derivatives, making it impractical to numerically perform such computations on the fly at each time step of our simulations. Thus, we precompute the auxiliary functions $\mathcal{I}_{k,\alpha}$ and $\mathcal{J}_{k,\alpha}$, that are directly related to the diffusion coefficients, for some logarithmically spaced values of $\mathcal{E}$ and $\psi$, and then, during the execution, we bilinearly interpolate these values of $\mathcal{I}_{k,\alpha}$ and $\mathcal{J}_{k,\alpha}$. In this way, we get approximate local values for those coefficients given the exact position and velocity of the stellar-mass BH at a particular time step.
@@
In these equations, $\Delta t$ is the length of the particular time step we are considering. Thus, the values of $\Delta v_{\parallel}$ and $\Delta v_{\perp}$ we draw represent a statistical realisation of the change in velocity due to all of the encounters that have occurred during that time step.
@@ 3.3 Orbit-averaged treatment of two-body relaxation
Figure 6: Sketch depicting an example orbit change following our orbit-averaged procedure. The stellar-mass BH, initially on the orbit labelled 00, is kicked when $r=a$, and ends on the orbit labelled $1$. Both orbits have the same semi-major axis $a$, but different eccentricities. The dashed circle represents a circular orbit with radius $a$. In this example there is no way to rotate either orbit around the focus occupied by the MBH such that the pericentre or apocentre of orbit 0 crosses orbit 1 at any point. Thus, neither pericentres nor apocentres are good spots to employ our orbit-averaged procedure. We note that the eccentricities of these example orbits are much lower compared to the ones investigated in this work. Moreover, the difference between subsequent orbits is much exaggerated here.
Figure 7: Evolution of the position of a stellar-mass BH over a few orbits with the orbit-averaged treatment of two-body relaxation. The top right panel shows the 3D trajectory of the stellar-mass BH, while the other three panels show its projections onto the three planes of the reference system. The colour of the curve shows the time elapsed since the beginning of the evolution. The red dots indicate the points where a velocity kick is given, close to $r=a$. We let the body complete the first orbit without applying kicks to visualise its trajectory better. The black cross shows the initial position of the stellar-mass BH, while the MBH is fixed at $(0,0,0)$. We note that the $y$- and $z$-axes have different scales for better readability. In this example $M_{\bullet}=4\times 10^{6}\,\mathrm{M}_{\sun}$, $a_{0}=0.1\,\mathrm{pc}$, $e_{0}=0.997$.
@@
In Appendix B we show that, once $\delta v_{\parallel}$ and $\delta v_{\perp}$ are drawn, then
@@
In Appendix C we show that a suitable orbit-integrated kick in velocity can be computed from $\Delta E$ and $\Delta J$ as
@@
When updating the orbit, we change the velocity of the stellar-mass BH without displacing it. This is possible only if the old and the new orbit intersect at the location where $\Delta v$ is applied, a fact that we ensure by applying the kick when $r\simeq a$ (Fig. 6 illustrates the situation with a sketch). It is safe to assume that the orbits intersect there, as long as the energy of the particle (and thus $a$, since $\mathcal{E}\simeq GM_{\bullet}/2a$) does not change drastically along a single period. [^10] We note that the $\Delta\mathcal{E}$ and $\Delta J$ we are using here only come from the effects of two-body relaxation. GWs emission can strongly and quickly modify the energy of the orbit, but their effect is only included in the 2.5PN term of the Hamiltonian. In case this assumption is not verified, the orbit-averaging approximation is not valid any more, as the time needed for two-body relaxation to significantly change $\mathcal{E}$ is less than the orbital period (Binney & Tremaine 2008; Merritt 2013a).
@@ 3.4 Stopping criteria and initial conditions
As our aim was to gather a statistically significant number of EMRIs and DPs, we chose to limit the parameter space that we explored to a sub-portion close to the loss cone edge. Moreover, as we assume a steady-state stellar profile, we are interested in exploring orbital parameters for which the relaxation time $t_{\mathrm{rlx}}$ is not too large, as the initial conditions and the diffusion coefficients would be affected by the overall relaxation of the system. In particular, we consider as our timescale of investigation ${\sim}1\%$ of the time-to-peak of (classical) EMRI rates in the systems we are considering, which estimates the timescale of evolution of the distribution function of scatterers (Broggi et al. 2022)
@@
where we leverage on the M–$\sigma$ relation (Gültekin et al. 2009) to obtain the final scaling with the MBH mass. We note that $t_{\mathrm{max}}$ is not to be intended as a dynamical stopping condition (meaning the simulations do not halt at the time $t=t_{\mathrm{max}}$), rather the run is ended only when the orbital elements cross over the $t_{\mathrm{rlx}}=t_{\mathrm{max}}$ curve on the $(1-e,a)$ plane.
We initialise all our simulations inside the region of interest where $t_{\mathrm{rlx}}<t_{\mathrm{max}}$ following a procedure described at the end of this section and we stop and discard any run where the orbit leaves the region. This choice does not affect the EMRI-to-DP ratio. In fact, $t_{\mathrm{max}}$ is so long with respect to the orbital period that any plunge or EMRI must enter this region before the capture. As long as the procedure is not biased towards EMRI or DP candidates, once a particle is discarded we are safe to reinitialise it as an uncorrelated particle.
The reinitialisation procedure becomes biased when the restriction to $t_{\mathrm{rlx}}<t_{\mathrm{max}}$ excludes a significant part of the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ curve. In this case, it can happen that some orbits well inside the GW emission dominated region but still far from the loss cone, can reach the $t_{\mathrm{rlx}}=t_{\mathrm{max}}$ boundary before crossing the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ separatrix, thus being erroneously stopped and reinitialised instead of being counted as EMRIs. A clear example of such an orbit is shown by the blue line in Fig. 8: the event will clearly produce an EMRI, but according to our reinitialisation criterion is effectively discarded and resimulated.
Figure 8: Two examples of orbit evolution in the $(1-e,a)$ plane. The blue line shows the formation of an EMRI that is missed by our standard definition, and requires the correction described in the text to be rightfully counted. The pink line instead shows a regular reinitialisation. We also plot the evolution line of an orbit that passes through the minimum of the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ curve following Eq. (32) (dashed brown line). It can be seen that all orbits that cross into the excluded region while being in the GW-dominated region fall below this evolution line, meaning they will reach the solid green curve, as long as the stochastic kicks due to two-body relaxation are small (see Fig. 1 for the meaning of the other elements). Here we set $M_{\bullet}=4\times 10^{6}\,\mathrm{M}_{\sun}$.
@@
As DP candidates are not affected by this issue, the EMRI-to-DP ratio is influenced by these discarded EMRIs. This problem mostly affects the EMRI-to-DP ratio at small initial semi-major axes for large $M_{\bullet}$, as we show in Sect. 4.1.
For this reason, we correct our counts of EMRIs to also include orbits which cross into the excluded region while inside the GW-dominated region. Before accepting the orbit as an EMRI, we also verify that the final value of $a$ is below the evolution line of an orbit which evolves as Eq. (32), where $c_{0}$ is chosen so that the line crosses the minimum of the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ curve (again, see Fig. 8). This means that, ignoring stochastic kicks due to two-body encounters, the orbit would eventually reach our condition for EMRI formation, while inside the excluded region. We also point out that near the $t_{\mathrm{rlx}}=t_{\mathrm{max}}$ curve, the condition $t_{\mathrm{GW}}<t_{\mathrm{rlx}}$ is likely sufficient to identify EMRIs anyway, as the stricter condition $t_{\mathrm{GW}}<10^{-3}\,t_{\mathrm{rlx}}$ is mainly introduced to avoid misclassifying DPs as EMRIs, and such a misclassification can only occur for $1-e\lesssim 10^{-4}$, where the $t_{\mathrm{GW}}=t_{\mathrm{rlx}}$ curve closely approaches the loss cone.
@@
at the corresponding energy $\mathcal{E}_{0}=GM_{\bullet}/2a_{0}$. Then, we randomly draw an initial value for $\mathcal{R}_{0}$ from the probability distribution function
@@
In our procedure, we initialise all particles at the apocentre of their initial orbits, rather than assigning a random location along the orbit. We select this point because PN corrections are least significant there and this choice does not bias the likelihood of a particle becoming a DP or an EMRI. Before reaching a disruptive pericentre, the particle will either quickly modify its orbital parameters through two-body scattering (when $t_{\mathrm{rlx}}\lesssim P$), effectively erasing any memory of the initial condition, or complete more than one period on the same orbit (when $t_{\mathrm{rlx}}\gtrsim P$), rendering the specific initial location along the orbit irrelevant.
@@ 4 Results
In this section we present the results we obtained running a large number of simulations with the code described in Sect. 3.
@@ 4.1 EMRI-to-DP ratio and cliffhanger EMRIs
Figure 10: Distribution of the maximum ratio between $a$ evaluated at subsequent apocentres for EMRIs for selected values of $M_{\bullet}$ and $a_{0}$. The black dotted vertical line shows $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}=2$, which roughly delimits cliffhanger EMRIs to its right. In the left panel, the orange dashed and green dash-dotted vertical lines show $a_{0}/a_{\mathrm{f}}^{\mathrm{lc}}$ from Eq. (38), which is a rough prediction of the highest possible $(a_{\alpha}/a_{\beta})_{\mathrm{max}}^{\mathrm{EMRI}}$ ratio. Each histogram is normalised to 1. We note the different $x$-axis scales and binning strategies in the two panels.
@@
We report our main findings in the left panel of Fig. 9. For $M_{\bullet}=10^{6}\,\mathrm{M}_{\sun}$ and $M_{\bullet}=4\times 10^{6}\,\mathrm{M}_{\sun}$, we recover the expected dichotomy between EMRIs and DPs. In this case, EMRIs can only form from orbits with $a_{0}\lesssim a_{\mathrm{c}}$, while those with $a_{0}\gtrsim a_{\mathrm{c}}$ exclusively result in DPs. We find that the transition between the two regimes occurs over about an order of magnitude in $a_{0}$ at $a_{\mathrm{c}}\simeq 10^{-2}R_{\mathrm{inf}}$, which is consistent within a factor of two with previous works (Hopman & Alexander 2005; Raveh & Perets 2021, QS24). We believe that the quantitative difference is driven by differences in the nuclear cluster models, in the accuracy of the dynamical model (e.g. PN order) and in the treatment of two-body relaxation. We better explore the impact that some of these factors have on the shape of $S(a_{0})$ in Sect. 4.2 and 4.3.
@@
Conversely, for $M_{\bullet}<10^{6}\,\mathrm{M}_{\sun}$, we observe a relevant number of EMRIs formed from initially wide orbits, with $a_{0}\gg a_{\mathrm{c}}$. These events are consistent with the cliffhanger EMRI population first seen in Monte Carlo simulations by QS24 and, to the best of our knowledge, never found in previous work. We observe cliffhanger EMRIs forming for $M_{\bullet}\lesssim 3\times 10^{5}\,\mathrm{M}_{\sun}$ and they appear to be more common as $M_{\bullet}$ decreases, consistently with the argument based on the $(1-e,a)$ phase space presented in Sect. 2.4.
@@
Our results for $S(a_{0})$ are compared to those obtained by QS24 and other authors in the right panel of Fig. 9. First, we can notice the large scatter in $a_{\mathrm{c}}/R_{\mathrm{inf}}$ that spans almost an order of magnitude, from $\approx 3\times 10^{-3}$ (QS24) to $\approx 2\times 10^{-2}$ (Raveh & Perets 2021). This is likely due to the differences in the parameters of the simulated systems as well as in the treatment of the dynamics and loss cone definition, as we discuss in Sect. 4.3. Focusing on cliffhanger EMRIs, QS24 observe their formation only for MBHs with $M_{\bullet}\lesssim 5\times 10^{4}\,\mathrm{M}_{\sun}$, a threshold set almost an order of magnitude below our limiting value. Moreover, we see a much larger fraction of EMRIs over DPs at large $a_{0}$, with a maximum ratio of $S(a_{0})\simeq 0.65$ at $a_{0}\gg 10^{-2}R_{\mathrm{inf}}$ for $M_{\bullet}=10^{4}\,\mathrm{M}_{\sun}$. As a further difference, our $S(a_{0})$ functions hint at a downward trend for $a_{0}\gg 10^{-2}R_{\mathrm{inf}}$, as opposed to tending to a constant at $a_{0}\to\infty$. This could be due to the fact that QS24 stopped their investigation at $a_{\mathrm{max}}\simeq R_{\mathrm{inf}}/3$, as their implementation does not account for the stellar potential.